Use the adder subtraction of Fig. 4.7 that has the following values of input Mode (M) and data inputs (A) and (B). In each case, determine the values of the outputs S3 S2 S1 S0 and C4.

            M            A                 B

            0          0111                0110

            1          101                  1010

            1         0000                 0001

The adder-subtractor circuit of Fig. 4-7 has the following values for input mode M and data inputs A and B. In each case, determine the values of the outputs: S3 S2 S1 S0 and C4.

M

A

B

a.

0

0111

0110

b.

1

0101

1010

c.

1

1000

0001
Answer #1

a) Mode is 0. 4-bit addition operation is performed. So, A(0111) + B(0110) gives 1101. where S3 = 1, S2 = 1, S1 = 0, S0 = 1 and C4 = 0 because there is no carry.

b) Mode is given as 1. So, 4-bit subtraction operation is performed on A and B.

A(0101) - B(1010) = 1) 1011 . where S3 = 1, S2 = 0, S1 = 1, S0 = 1 and C4 = 1 because there is 1 bit borrow. Result in decimal form is -5.

c) Mode is given as 1. So, 4-bit subtraction operation is performed on A and B.

A(1000) - B(0001) = 0111 . where S3 = 0, S2 = 1, S1 = 1, S0 = 1 and C4 = 0 because there is no borrow.